OpenStudy (anonymous):

(((((((MEDAL AND FAN))))))))))) Write an indirect proof proving the two non-right angles of a right triangle measure less than 90°.

OpenStudy (anonymous):

sum of all angles is 180 a+b+c= 180 and one of the angle let say c is 90 then a+b+90 = 180 which gives a+b = 90 a>0 and b>0 and there is addition is 90 .. what does that mean .. that means both are less than 90

OpenStudy (anonymous):

oh okay, I think I'm getting it. But would that be an indirect proof?

OpenStudy (anonymous):

ok let proof indirectly

OpenStudy (anonymous):

see in right angle triangle. one angle is 90. if another angle is also bigger than 90 then sum will become more than 180 .. which is not possible

OpenStudy (anonymous):

did you understand

OpenStudy (anonymous):

Okay Ive got that part

OpenStudy (anonymous):

So in the proof above, A and B are both less than 90?